OpenStudy (anonymous):

Given f(2)=4 and f'(2)=-5 find the equation of the tangent line at x=2

OpenStudy (bahrom7893):

y - 2 = -5(x-4)

OpenStudy (bahrom7893):

Slope is the tangent line which is the first derivative..

OpenStudy (anonymous):

shouldnt you switch the 2 and the 4?

OpenStudy (anonymous):

please show working lol thanks

OpenStudy (bahrom7893):

oh yea sorry...

OpenStudy (anonymous):

or am i tired?

myininaya (myininaya):

y=mx+b m=f'(2)=-5 y=-5x+b x=2 y=4 4=-5*2+b 14=b so y=-5x+14

OpenStudy (bahrom7893):

you're right joemath lol i just woke up haha..

OpenStudy (anonymous):

because your crazy smart you made me second guess myself lolol

OpenStudy (anonymous):

thanks joe! myininaya and bahroom!

OpenStudy (bahrom7893):

Here's the deal purple: f ' (x) = slope f ( x ) = y So you have a point: (x,y); in your case (2,4) slope, m = f ' (2) = -5 y-y1 = M(x-x1) y-4=-5(x-2)

OpenStudy (anonymous):

and bahrom* go get some cofee thanks

OpenStudy (bahrom7893):

lol haha.. i will..

OpenStudy (anonymous):

LOL

myininaya (myininaya):

i want a java chip frapaccino so bad!

OpenStudy (anonymous):

lol...myinaya i could use a caramel frap from starbucks

myininaya (myininaya):

you need to try the java chip

OpenStudy (anonymous):

i can never remember the name of what i get, i add peppermint to it though <.<

OpenStudy (anonymous):

lol...java chip probably isnt in my area im out guys thanks for your help!

myininaya (myininaya):

later